# Cross-Modality Perturbation Synergy Attack for Person Re-identification

### v2

@@ 3.0.1 Overall Framework
3.0.2 Optimizing Loss Functions for Attacking
In this study, our objective is to deceive cross-modality ReID models using a universal perturbation. Inspired by Bouniot et al. (2020), we employ a method utilizing triplet loss to bring similar pairs closer and push dissimilar pairs apart. Unlike Bouniot et al. (2020), in our triplet loss approach, we not only leverage visible image data but also introduce non-visible images (including grayscale images) to influence the distance relationships between images from different modalities.
We follow the approach of Li et al. (2019) to optimize the perturbation using cluster centroids. This method directly impacts the similarity between pedestrian identities in the ReID model’s feature space (rather than the similarity between individual samples), making it more effective. Subsequently, leveraging the acquired cluster centroids, we apply triplet loss to distort the pairwise relations between pedestrian identities. This process can be represented as follows:
[…]
As shown in Fig. 3, the loss function mentioned above fully leverages the triplet-wise relationships across different modality. Through this loss, we are able to pull the negative samples of each modality closer to the adversarial samples and push the positive samples of each modality away from the adversarial samples. Here, $C^{p}_{v}$ and $C^{n}_{v}$ represent the cluster centroids of the positive samples to push and negative samples to pull, respectively, in the original visible image feature space of the training data. Similar definitions apply to other modalities. $f^{adv}_{v}$, $f^{adv}_{g}$, and $f^{adv}_{ir}$ denote the perturbed features of the disturbed image in the visible, grayscale, and infrared (or thermal) modalities, respectively.
Figure 4: Schematic diagram of cross-modality attack augmentation.
3.0.3 Cross-Modality Attack Augmentation Method
Data augmentation stands as a crucial technique in deep learning, involving purposeful transformations or expansions of training data to enhance the performance and generalization of deep learning models. It is almost indispensable in deep learning tasks due to its effectiveness in increasing the diversity of training data, aiding models in better capturing the complexity of the data.
Intuitively, as depicted in Fig. 5, if data from different modalities encompass as many common factors as possible, the learned perturbation will more easily capture their shared features. Homogeneous grayscale images serve as effective bridges between different modalities in this context.
Drawing inspiration from Gong et al. (2021), we integrate random grayscale transformations into adversarial attack methods, termed Cross-Modality Attack Augmentation. This procedure involves guiding cross-modality perturbations by utilizing homogeneous grayscale images obtained from samples of various modalities. The specific objective is to explore structural relationships across diverse modalities.
@@ 3.0.3 Cross-Modality Attack Augmentation Method
$x_{g}=[t(R,G,B),t(R,G,B),t(R,G,B)],$ (3)
3.0.4 Cross-Modality Perturbation Synergy Attack
To synergistically leverage gradient information from diverse modalities for perturbation optimization, narrow the gap between different modalities to better capture shared features, we employ a three-step training process for the universal perturbation.
(1) Learning the visible modality. We employ visible images to learn the perturbation $\eta$ using the loss functions introduced in Eq. 1. For a given batch of visible images with $n$ samples, we perturb and extract their features using the cross-modality ReID model. The corresponding loss function for this step is formulated as:
$L_{RGB}(F^{adv}_{RGB},\eta)=\frac{1}{n}\sum_{i=1}^{n}[L_{tri}(f^{adv}_{RGB},\eta)],$ (4)
where $F^{adv}_{RGB}$ represents $n$ features of the current disturbed visible image batch, and $f^{adv}_{RGB}$ denotes the $i$-th perturbed feature. We adopt the approach proposed in Li et al. (2019) to iteratively update the temporary perturbation $\eta$ using Momentum-Inertia Stochastic Gradient Descent (MI-SGD), expressed as:
@@ 3.0.4 Cross-Modality Perturbation Synergy Attack
Here, $\theta$ represents the momentum value (set as $\theta=1$), and $\Delta^{\prime}_{ir}$ derived from the previous iteration. The iteration step size is denoted by $\alpha$ (set as $\alpha=\frac{\epsilon}{12}$ ), where $\epsilon$ is the adversarial bound ($\epsilon=8$, unless otherwise specified). We set the margin $\rho=0.5$ in triplet loss.
(2) Learning the grayscale modality. This part is executed through our data augmentation. It is not considered as a separate module and is therefore not explicitly listed in Alg. 1.
Specifically, during the perturbation learning process, we randomly transform visible or infrared (or thermal) images into homogeneous grayscale images, participating in the iterative optimization of adversarial perturbations. It is employed to bridge the gap between different modalities, thereby improving the universality of the perturbation across diverse modalities.
(3) Learning the infrared modality. This step is similar to (1); We utilize the infrared (or thermal) images to learn the perturbation $\eta$ with the loss functions:
$L_{ir}(F^{adv}_{ir},\eta)=\frac{1}{n}\sum_{i=1}^{n}[L_{tri}(f^{adv}_{ir},\eta)]$ (8)
@@
The main difference compared to the previous step lies in the perturbation applied to the input and the gradients related to momentum.
@@ 4 Experiments
In this section, we compare our approach with several methods, including M-FGSM Bai et al. (2020), as well as state-of-the-art methods such as LTA Gong et al. (2022), Col.+Del. Yang et al. (2023) and ODFA Zheng et al. (2023).
Table 1: Results for attacking cross-modality ReID systems on the SYSU dataset. It reports on visible images querying infrared images and vice versa. Rank at $r$ accuracy (%) and mAP (%) are reported.
@@
AGW baseline TPAMI 2022 47.50 84.39 92.14 47.65 54.17 91.14 95.98 62.97
M-FGSM attack TPAMI 2020 25.79 49.04 57.96 19.24 20.56 38.91 46.35 15.89
ODFA attack IJCV 2023 25.43 47.49 56.38 19.00 14.62 29.92 36.42 11.35
ours attack ———— 1.11 8.67 16.14 1.41 1.31 7.47 10.36 1.23
DDAG baseline ECCV 2020 54.75 90.39 95.81 53.02 61.02 94.06 98.41 67.98
M-FGSM attack TPAMI 2020 28.36 52.47 60.76 23.11 24.85 40.74 49.22 18.40
ODFA attack IJCV 2023 27.75 50.26 59.14 22.30 17.62 32.64 40.03 14.83
ours attack ———— 1.62 7.59 14.46 1.84 1.45 7.71 10.72 1.25
Table 2: Results for attacking cross-modality ReID systems on the RegDB dataset. It reports on visible images querying thermal images and vice versa. Rank at $r$ accuracy (%) and mAP (%) are reported.
@@
AGW baseline TPAMI 2022 70.05 86.21 91.55 66.37 70.49 87.21 91.84 65.90
ours attack ———— 2.29 9.06 18.35 3.92 1.93 11.44 19.30 3.46
DDAG baseline ECCV 2020 69.34 86.19 91.49 63.46 68.06 85.15 90.31 61.80
ours attack ———— 1.33 10.28 19.06 3.79 1.35 9.52 17.52 3.19
Table 3: Transferability experiments of the proposed method across different models (visible to thermal).
Settings RegDB
Method State $r=1$ $r=10$ $r=20$ mAP
AGW Before attack 70.05 86.21 91.55 66.37
AGW→DDAG After attack 3.91 11.84 17.45 5.27
DDAG Before attack 69.34 86.19 91.49 63.46
DDAG→AGW After attack 3.02 10.84 17.47 4.12
Datasets. We evaluate our proposed method on two commonly used cross-modality ReID datasets: SYSU-MM01 Wu et al. (2017) and RegDB Nguyen et al. (2017). SYSU-MM01 is a large-scale dataset with 395 training identities, captured by 6 cameras (4 RGB, 2 near-infrared) on the SYSU campus. It comprises 22,258 visible and 11,909 near-infrared images. The testing set consists of 95 identities with two evaluation settings. The query sets include 3803 images from two IR cameras. We conduct ten trials following established methods Wang et al. (2019); Ye et al. (2020b) and report the average retrieval performance. Please refer to Wu et al. (2017) for the evaluation protocol.
RegDB Nguyen et al. (2017) is a smaller-scale dataset with 412 identities, each having ten visible and ten thermal images. we randomly select 206 identities (2,060 images) for training and use the remaining 206 identities (2,060 images) for testing. We assess our model in two retrieval scenarios: visible-thermal and thermal-visible performance.
Evaluation Metrics. Following existing works Zheng et al. (2015), we employ Rank-k precision and Cumulative Matching Characteristics (CMC) and mean Average Precision (mAP) as evaluation metrics. Rank-1 represents the average accuracy of the top-ranked result corresponding to each cross-modality query image. mAP represents the mean average accuracy, where the query results are sorted based on similarity, and the closer the correct result is to the top of the list, the higher the precision. Please note that, for adversarial attacks, a lower accuracy indicates a more successful attack.
4.0.1 Performance on Cross-Modality ReID
We used AGW Ye et al. (2022) and DDAG Ye et al. (2020a) as baseline models for testing on the RegDB and SYSU cross-modality ReID datasets.
AGW (Attention Generalized mean pooling with Weighted triplet loss) enhances the learning capability of crucial features by integrating non-local attention blocks, learnable GeM pooling, and weighted regularization triplet loss.
DDAG (Dynamic Dual-Attentive Aggregation) improves feature learning by combining intra-modality weighted-part attention and cross-modality graph structured attention, considering both part-level and cross-modal contextual cues.
@@ 4.0.1 Performance on Cross-Modality ReID
The experiments encompass two scenarios: 1) Perturbing visible images (query) to disrupt the retrieval of infrared or thermal non-visible images (gallery). This is denoted as ”Visible to Infrared” in Tab.1 and ”Visible to Thermal” in Tab.2. 2) Perturbing infrared or thermal non-visible images (query) to interfere with the retrieval of visible images (gallery). This is indicated as ”Infrared to Visible” in Tab.1 and ”Thermal to Visible” in Tab.2.
@@
From Table 1, it can be observed that the proposed method, whether in ”Visible to Infrared” or ”Infrared to Visible” causes the model accuracy to drop below 2% on the rank-1. Similarly, from Table 2, it can be observed that the proposed method, whether in ”Visible to Thermal” or ”Thermal to Visible” leads to a model accuracy drop below 3% on the rank-1. In contrast, traditional attack methods Metric-FGSM (M-FGSM) and ODFA result in attacked models with accuracy above 14% on rank-1. These results indicate: 1) In comparison to methods that only utilize single-modality information, our proposed method has a significant attacking effect on different modalities. 2) From the perspective of attack effectiveness, cross-modality retrieval such as ”Thermal to Visible” appears to be relatively less susceptible to attacks (higher mAP than ”Infrared to Visible”). We speculate that this may be related to thermal having less texture information, reducing the discernible features for attackers and consequently lowering the success rate of attacks.
Table 4: Transferability experiments of the proposed method across different models (visible to Infrared).
Settings SYSU
Method State $r=1$ $r=10$ $r=20$ mAP
AGW Before attack 47.50 84.39 92.14 47.65
AGW→DDAG After attack 2.07 10.39 19.35 3.41
DDAG Before attack 54.75 90.39 95.81 53.02
DDAG→AGW After attack 1.94 9.46 17 .53 2.80
4.0.2 Transferability of CMPS
@@
From Tab.3 and Tab.4, the results of the proposed method’s transfer attacks on two baseline models, AGW and DDAG, can be observed. For example, on the SYSU dataset, the original attack result of the proposed method on DDAG is mAP=1.41% (refer to Tab. 1). When the perturbation is transferred from AGW to DDAG, the attack result becomes mAP=3.41%. This indicates that the proposed attack method exhibits good generalization across different models, and thus, the attack performance does not degrade significantly. This consistent result is observed on both the RegDB and SYSU datasets.
4.0.3 Ablation Study

### v3

@@ 3.3 Cross-Modality Attack Augmentation Method
Figure 4: Cross-modality attack augmentation: bridging gap between visible and non-visible (infrared) modalities with grayscale.
@@
$x_{g}=[t(R,G,B),t(R,G,B),t(R,G,B)].$ (3)
3.4 Cross-Modality Perturbation Synergy Attack
To synergistically utilize gradient information from diverse modalities for perturbation optimization, narrow the gap between different modalities to better capture shared knowledge, we adopt the following training process to generate a universal perturbation:
(1) Learning the visible modality. For a given batch of visible images with $n$ samples, we extract and perturb their features using the cross-modality ReID model. We update the temporary perturbation $\eta$ iteratively using Momentum-Inertia Stochastic Gradient Descent (MI-SGD), expressed as:
@@ 3.4 Cross-Modality Perturbation Synergy Attack
Here, $\theta$ represents the momentum value (set as $\theta=1$), and $\Delta^{\prime}_{ir}$ is derived from the previous iteration. The iteration step size is denoted by $\alpha$ (set as $\alpha=\frac{\epsilon}{12}$ ), where $\epsilon$ is the adversarial bound ($\epsilon=8$, unless otherwise specified). We set the margin $\rho=0.5$ in our triplet loss.
(2) Learning the grayscale modality. This part is executed through data augmentation. It is not considered as a separate module and is therefore not explicitly listed in Alg. 1. Specifically, during the perturbation learning process, we randomly transform visible or infrared (or thermal) images into homogeneous grayscale images, participating in the iterative optimization of adversarial perturbations. It is employed to bridge the gap between different modalities, thereby improving the universality of the perturbation across diverse modalities. In order to investigate the impact of different grayscale conversion probabilities on attack performance, we conducted a series of ablation experiments. For details, please refer to Fig. 5 in Supplementary Material.
(3) Learning the infrared (or thermal) modality. This step is similar to (1). We utilize the infrared (or thermal) images to learn the perturbation $\eta$ with the our loss functions:
@@
Here, $\Delta_{RGB}$ derived from step (1). The main difference compared to the previous step lies in the perturbation applied to the input and the gradients related to momentum.
Theoretical Analysis. In traditional optimization, after optimizing for one modality, the perturbation may become suboptimal for the other, resulting in higher errors and bias towards a single modality. In contrast, the proposed aggregated optimization method jointly optimizes both modalities by simultaneously considering their losses, ultimately finding a universal perturbation that improves cross-modality performance. In the supplementary material 7, we provide a mathematical analysis demonstrating the effectiveness of this method compared to traditional attack methods that lack intrinsic correlations between different modalities.
@@ 4 Experiments
In this section, we compare our approach with several methods, including traditional classification attack methods FGSM [11] and PGD [19], traditional metric attack methods like Metric-FGSM [1], as well as state-of-the-art ReID attack methods such as LTA [6] The LTA code is available at: https://github.com/finger-monkey/LTA_and_joint-defence , ODFA[47] and Col.+Del.[39]. Our experiments were conducted using three RTX 2080 Ti GPUs, each with 11GB of memory.
Datasets. We evaluate our proposed method on two commonly used cross-modality ReID datasets: SYSU-MM01 [34], RegDB [22] and LLCM [43]. SYSU-MM01 is a large-scale dataset with 395 training identities, captured by 6 cameras (4 RGB, 2 near-infrared) on the SYSU campus. It comprises 22,258 visible and 11,909 near-infrared images. The testing set consists of 95 identities with two evaluation settings. The query sets include 3803 images from two IR cameras. We conduct ten trials following established methods [33] and report the average retrieval performance. Please refer to [34] for the evaluation protocol. RegDB [22] is a smaller-scale dataset with 412 identities, each having ten visible and ten thermal images. we randomly select 206 identities (2,060 images) for training and use the remaining 206 identities (2,060 images) for testing. We assess our model in two retrieval scenarios: visible-thermal and thermal-visible performance. LLCM is a dataset designed specifically for cross-modality ReID in low-light environments. Compared to other datasets, its diverse scenarios and low-light conditions present greater challenges for attackers. This complexity and uncertainty make adversarial attacks more difficult to execute.
Evaluation Metrics. Following existing works [45], we employ Rank-k precision and Cumulative Matching Characteristics (CMC) and mean Average Precision (mAP) as evaluation metrics. Rank-1 represents the average accuracy of the top-ranked result corresponding to each cross-modality query image. mAP represents the mean average accuracy, where the query results are sorted based on similarity, and the closer the correct result is to the top of the list, the higher the precision. Please note that, for adversarial attacks, a lower accuracy indicates a more successful attack.
Table 1: Results for attacking cross-modality ReID systems on the SYSU [34] dataset. It reports on visible images querying infrared images and vice versa. Rank at $r$ accuracy (%) and mAP (%) are reported. For the "Visible to Infrared" scenario, we used the all-search mode. For the "Infrared to Visible" scenario, we used the indoor-search mode.
@@
AGW baseline [41] TPAMI 2022 47.50 84.39 92.14 47.65 54.17 91.14 95.98 62.97
FGSM attack [11] ICLR 2015 42.64 81.21 89.32 43.67 48.05 86.73 92.11 53.22
PGD attack [19] ICLR 2018 39.14 76.80 85.42 40.91 43.68 82.54 89.14 48.56
M-FGSM attack [1] TPAMI 2020 25.79 49.04 57.96 19.24 20.56 38.91 46.35 15.89
LTA attack [6] CVPR 2022 8.42 21.25 27.98 9.16 20.92 32.18 36.80 15.24
ODFA attack [47] IJCV 2023 25.43 47.49 56.38 19.00 14.62 29.92 36.42 11.35
Col.+Del. attack [39] TPAMI 2023 3.23 14.48 20.15 3.27 4.12 16.85 21.27 3.89
Our attack   1.11 8.67 16.14 1.41 1.31 7.47 10.36 1.23
DDAG baseline [40] ECCV 2020 54.75 90.39 95.81 53.02 61.02 94.06 98.41 67.98
FGSM attack [11] ICLR 2015 48.27 86.02 91.34 49.55 53.87 90.15 94.58 57.84
PGD attack [19] ICLR 2018 50.62 88.30 93.12 51.89 56.10 91.54 96.13 59.22
M-FGSM attack [1] TPAMI 2020 28.36 52.47 60.76 23.11 24.85 40.74 49.23 18.40
LTA attack [6] CVPR 2022 10.54 23.08 30.47 12.28 18.93 34.12 41.52 15.04
ODFA attack [47] IJCV 2023 27.75 50.26 59.14 22.30 17.62 32.64 40.03 14.83
Col.+Del. attack [39] TPAMI 2023 4.28 16.12 21.36 3.97 6.28 19.53 25.61 5.21
Our attack NeurIPS 2024 1.62 7.59 14.46 1.84 1.45 7.71 10.72 1.25
Table 2: Results for attacking cross-modality ReID systems on the RegDB [22] dataset. It reports on visible images querying thermal images and vice versa. Rank at $r$ accuracy (%) and mAP (%) are reported.
@@
AGW baseline [41] TPAMI 2022 70.05 86.21 91.55 66.37 70.49 87.21 91.84 65.90
FGSM attack [11] ICLR 2015 66.79 83.14 88.46 61.05 65.42 81.98 87.20 60.12
PGD attack [19] ICLR 2018 62.14 80.28 85.10 57.34 63.71 78.82 84.05 58.42
M-FGSM attack [1] TPAMI 2020 29.34 52.90 61.44 23.35 23.64 40.36 48.61 18.57
LTA attack [6] CVPR 2022 12.65 25.24 34.02 12.80 10.51 22.93 31.79 9.74
ODFA attack [47] IJCV 2023 28.57 51.42 60.58 21.84 17.26 33.27 42.92 15.27
Col.+Del. attack [39] TPAMI 2023 5.12 16.83 22.10 4.94 4.92 14.47 23.04 4.86
Our attack   2.29 9.06 18.35 3.92 1.93 11.44 19.30 3.46
DDAG baseline [40] ECCV 2020 69.34 86.19 91.49 63.46 68.06 85.15 90.31 61.80
FGSM attack [11] ICLR 2015 61.83 80.12 86.47 55.78 60.94 78.35 84.09 56.91
PGD attack [19] ICLR 2018 64.58 81.39 87.20 58.45 62.17 79.02 85.27 57.69
M-FGSM attack [1] TPAMI 2020 30.86 54.16 61.98 24.01 25.83 42.12 49.76 19.33
LTA attack [6] CVPR 2022 11.65 23.20 32.73 11.41 9.76 21.53 29.96 9.23
ODFA attack [47] IJCV 2023 29.64 52.74 60.74 23.88 24.06 39.75 46.25 18.64
Col.+Del. attack [39] TPAMI 2023 4.68 13.55 18.57 4.39 4.23 12.75 20.82 4.05
Our attack NeurIPS 2024 1.33 10.28 19.06 3.79 1.35 9.52 17.52 3.19
4.1 Performance on Cross-Modality ReID
We used AGW [41] and DDAG [40] as baseline models for testing on the RegDB and SYSU cross-modality ReID datasets. AGW (Attention Generalized mean pooling with Weighted triplet loss) enhances the learning capability of crucial features by integrating non-local attention blocks, learnable GeM pooling, and weighted regularization triplet loss. DDAG (Dynamic Dual-Attentive Aggregation) improves feature learning by combining intra-modality weighted-part attention and cross-modality graph structured attention, considering both part-level and cross-modal contextual cues. Additionally, we use DEEN [43] (Diverse Embedding Expansion Network) as baseline models for testing on the LLCM [43] cross-modality ReID datasets. The core idea of DEEN is to enhance the feature representation capability by introducing a diversity embedding mechanism. The network expands the embedding space, allowing features from visible and infrared images to align better in a high-dimensional space, thereby improving the accuracy of cross-modality matching.
@@ 4.1 Performance on Cross-Modality ReID
The experiments encompass two scenarios: 1) Perturbing visible images (query) to disrupt the retrieval of infrared or thermal non-visible images (gallery). This is denoted as "Visible to Infrared" in Tab.1 and "Visible to Thermal" in Tab.2. 2) Perturbing infrared or thermal non-visible images (query) to interfere with the retrieval of visible images (gallery). This is indicated as "Infrared to Visible" in Tab.1 and "Thermal to Visible" in Tab.2.
@@
From Tab.1, it can be seen that the proposed method reduces the rank-1 accuracy to below 2% in both the ’Visible to Infrared’ and ’Infrared to Visible’ cases. Similarly, from Tab.2, the rank-1 accuracy drops below 3% in both the ’Visible to Thermal’ and ’Thermal to Visible’ scenarios. In contrast, traditional metric-based attacks like Metric-FGSM (M-FGSM)[1] and ODFA[47] result in attacked models with rank-1 accuracy above 14%, while traditional classification attacks (such as FGSM [11] and PGD [19]) perform even worse, with rank-1 accuracy remaining over 60%. This is because ReID relies on metric learning for feature matching rather than category classification, requiring attacks specifically tailored for metric learning. These results indicate: 1) Compared to traditional methods that optimize perturbations separately for each modality without considering the inherent correlations between different modalities, our proposed approach demonstrates significant attacking effectiveness across different modalities. 2) From the perspective of attack effectiveness, cross-modality retrieval scenarios such as ’Thermal to Visible’ appear to be relatively less susceptible to attacks (with higher mean Average Precision than ’Infrared to Visible’). We speculate that this may be due to thermal infrared images having fewer texture details, reducing discernible features for attackers and thus lowering the success rate of attacks.
Table 3: Results for attacking cross-modality ReID systems on the LLCM [43] dataset. It reports on visible images querying thermal images and vice versa. Rank at $r$ accuracy (%) and mAP (%) are reported.
Settings Visible to Infrared Infrared to Visible
Method Venue $r=1$ $r=10$ $r=20$ mAP $r=1$ $r=10$ $r=20$ mAP
DEEN baseline [41] CVPR 2023 62.53 90.31 94.73 65.84 54.96 84.92 90.91 62.95
M-FGSM attack [1] TPAMI 2020 28.48 64.92 75.12 32.88 25.64 61.45 78.31 30.46
LTA attack [6] CVPR 2022 15.16 56.42 67.53 21.47 19.54 58.25 70.72 24.86
ODFA attack [47] IJCV 2023 26.34 65.24 76.92 30.85 23.73 62.46 73.57 29.63
Col.+Del. attack [39] TPAMI 2023 8.61 22.73 36.07 15.72 9.13 20.76 38.02 16.31
Our attack NeurIPS 2024 5.83 18.14 27.56 12.47 6.42 19.53 28.54 12.23
Comparison with State-of-the-Art. Col.+Del., as a universal perturbation method, was fairly compared by first optimizing with one modality’s dataset and then fine-tuning with the other modality. Since universal perturbations capture shared patterns across the entire data distribution, Col.+Del. is capable of achieving some level of attack effectiveness in cross-modality scenarios. However, by comparing Tab.1, Tab.2, and Tab.3, we observe that although Col.+Del. performs better than other methods, its effectiveness is still noticeably limited due to the lack of intrinsic correlation mechanisms between modalities. Moreover, as shown in Fig.6, our method outperforms Col.+Del. in transfer attacks across different baselines in cross-modality ReID. The conclusions from these experiments are as follows: 1) In cross-modality attacks, Col.+Del. demonstrates the feasibility of universal perturbations. However, its performance is limited by its failure to account for modality differences and inherent correlations. 2) Our method better bridges the gap between different modalities, more effectively capturing shared features across them.
4.2 Transferability of CMPS
@@
From Fig.6 in supplemental material, the results of the proposed method’s transfer attacks on two baseline models, AGW and DDAG, can be observed. For example, on the SYSU dataset, the original attack result of the proposed method on DDAG is mAP=1.84% (refer to Tab. 1). When the perturbation is transferred from AGW to DDAG, the attack result becomes mAP=3.41%. This indicates that the proposed attack method exhibits good generalization across different models, and thus, the attack performance does not degrade significantly. This consistent result is observed on both the RegDB and SYSU datasets. Similarly, in Fig.7 of the supplemental material, we evaluate the cross-dataset transferability of perturbations in comparison with Col.+Del. The results demonstrate a significant advantage of our method. Additionally, we conducted adversarial transferability experiments on IDE [46], PCB [26], and ResNet18 [12]. The rank-1 transfer attack success rates are presented in Tab.4. It can be observed that our method consistently achieves higher transfer attack success rates across all model combinations compared to Col.+Del., indicating that our method demonstrates stronger robustness in generating more universal adversarial perturbations.
Table 4: Comparison of transfer attack success rates between our method and Col.+Del. across models, with higher values indicating better transferability.
[…]
IDE 98.7% / 94.3% 84.5% / 81.2% 87.4% / 86.1%
PCB 85.1% / 80.4% 97.6% / 92.8% 88.3% / 85.7%
ResNet18 81.0% / 78.5% 77.5% / 74.9% 98.2% / 95.6%
4.3 Ablation Study